Three point charges are positioned on the x-axis: +117.0 μC at x = 0.00 cm, +16.0 μC at x = 40.0 cm, and -62.0 μC at x = 50.0 cm. What is the magnitude of the electrostatic force acting on the +117.0–μC charge?
Answer #1

Here ,

electric force on the third charge is F

force between two charges is F = k * q1*q2/d^2

as F = F1 - F2

F = 117 *10^-6 * 9*10^9 * 10^-6 * (62/5^2 - 16/.4^2)

F = 155.8 N

the magnitude of electric force on the 117 uC charge is 155.8 N